# AutoCodeRover: Autonomous Program Improvement

### v2

@@ Abstract.
Researchers have made significant progress in automating the software development process in the past decades. Automated techniques for issue summarization, bug reproduction, fault localization, and program repair have been built to ease the workload of developers. Recent progress in Large Language Models (LLMs) has significantly impacted the development process, where developers can use LLM-based programming assistants to achieve automated coding. Nevertheless software engineering involves the process of program improvement apart from coding, specifically to enable software maintenance (e.g. program repair to fix bugs) and software evolution (e.g. feature additions). In this paper, we propose an automated approach for solving Github issues to autonomously achieve program improvement. In our approach called AutoCodeRover, LLMs are combined with sophisticated code search capabilities, ultimately leading to a program modification or patch. In contrast to recent LLM agent approaches from AI researchers and practitioners, our outlook is more software engineering oriented. We work on a program representation (abstract syntax tree) as opposed to viewing a software project as a mere collection of files. Our code search exploits the program structure in the form of classes/methods to enhance LLM’s understanding of the issue’s root cause, and effectively retrieve a context via iterative search. The use of spectrum based fault localization using tests, further sharpens the context, as long as a test-suite is available. We consider the recently proposed SWE-bench-lite which consists of 300 real-life Github issues involving bug fixing and feature additions. Experiments on SWE-bench-lite show increased efficacy in solving Github issues (22-23% on SWE-bench-lite), as compared to recent efforts from the AI community. Interestingly, our approach resolved 67 GitHub issues in less than 12 minutes each, whereas developers spent more than 2.77 days on average. On the full SWE-bench consisting of 2294 GitHub issues, AutoCodeRover solved around 16% of issues, which is higher than the efficacy of the recently reported AI software engineer Devin from Cognition Labs, while taking time comparable to Devin. We posit that our workflow enables autonomous software engineering, where, in future, auto-generated code from LLMs can be autonomously improved.
ccs: Software and its engineering Automatic programmingccs: Software and its engineering Software testing and debuggingccs: Computing methodologies Natural language processing
@@ 1. Beyond Automatic Programming
Our contribution lies in the effective use of code search to make software engineering processes like program repair autonomous. We demonstrate this capability by autonomously solving GitHub issues. We report favorable experimental results on the SWE-bench dataset (Jimenez et al., 2024b). We achieve 15.95% efficacy on the full SWE-bench with 2294 GitHub issues, and 22.33% efficacy on the SWE-bench-lite subset with 300 GitHub issues. Overall, we see the need to endow a software engineering oriented outlook to the recent flurry of activity on LLM agents for software engineering, which mostly has only an AI flavor. We believe that the software engineering angle can be conveyed via (at least) the following five dimensions, and can enhance the capabilities of LLM agents.
@@
  • We posit that higher efficacy of automated repair is more important than time efficiency, as long as the time is within a threshold. It is well-known in empirical software engineering research that time limits of 30-60 minutes for automated repair are tolerable, based on extensive developer surveys in the field (Noller et al., 2022). We thus report 22% efficacy on SWE-bench-lite in solving GitHub issues, within 12 minutes. We can compare this timelimit of 12 minutes to the average time of 2.77 days to fix the GitHub issues manually.
  • One should be able to exploit debugging techniques like test-based fault localization to guide the search for code in resolving GitHub issues. We show that use of fault localization in setting the code context leads to an increase in efficacy of AutoCodeRover in solving GitHub issues.
@@ 2.1. Program Repair
Recent work (Fan et al., 2023; Jiang et al., 2023; Xia et al., 2023) have shown the use of LLMs for automated program repair. This line of work often assumes the buggy program statements are given (i.e. perfect fault localization assumption) and focuses on constructing APR-specific prompts that guide LLM to generate a patch for the selected buggy program statements multiple times until a patch that passes all tests is found. However, obtaining buggy locations for a large project is an essential and challenging task in resolving real-life bug reports.
@@
2.2. Dataset
SWE-bench (Jimenez et al., 2024b) is a benchmark that aims to evaluate the capabilities of large language models in resolving end-to-end real-life software engineering tasks. The benchmark consists of 2,294 real-life software engineering task instances collected from the repositories of 12 popular large Python projects  [^1] https://www.swebench.com/index.html (e.g., django, sympy). Each SWE-bench task instance contains a pair of Github issue and corresponding pull requests. The GitHub issue either reports a bug to be fixed or requests to implement new features. The pull request includes the code changes made by human developers to resolve the issue and test cases that prevent the issue. Unlike traditional code generation tasks in HumanEval (Chen et al., 2021) and MBPP (Li et al., 2022) benchmark, resolving a SWE-bench instance is particularly challenging because it requires automatically generating code changes that address the problem in a GitHub issue for a matured large code repository based only on the issue description. More specifically, the process may involve a series of complex tasks like reasoning the target bug location across files in the code repository, analyzing the root cause of the issue, proposing bug-fixing strategies, and eventually writing a patch that passes all the test cases added in the pull request.
SWE-bench team also constructed a lite version of SWE-bench (SWE-bench lite) due to the high cost of evaluating in the complete SWE-bench. SWE-bench lite includes 300 task instances sampled from SWE-bench, following a similar repository distribution. SWE-bench lite removed task instances with meta files such as images, external links, and instances without proper description (e.g., short problem statement which is fewer than 40 words).
@@ 3. Motivating Example
1""" AutoCodeRover patch. """
2 except (ValueError, TypeError, self.queryset.model.DoesNotExist):
3- raise ValidationError(self.error_messages[’invalid_choice’],
4- code=’invalid_choice’)
5+ % Include the invalid value in the error message
6+ raise ValidationError(
7+ self.error_messages[’invalid_choice’] % {’value’: value},
8+ code=’invalid_choice’
9+ )
10
11""" Developer patch. """
12 except (ValueError, TypeError, self.queryset.model.DoesNotExist):
13- raise ValidationError(self.error_messages[’invalid_choice’],
14- code=’invalid_choice’)
15+ raise ValidationError(
16+ self.error_messages[’invalid_choice’],
17+ code=’invalid_choice’,
18+ params={’value’: value},
19+ )
@@
In Step 3, the agent receives the implementation of validate and to_python methods. Among the two methods, it selects to_python to be the more suitable place to make changes, since to_python raises the relevent exception and does not include the invalid value. At this point, the retrieval agent deems the collected code context as sufficient for understanding the issue and drafting the patch. The identified buggy location, together with the gathered context and analysis so far, is passed to another patch generation agent. This agent is instructed to write patches following the format described in Step 4 (see yellow box) of Figure 2. In Step 4, a patch is written to allow a value to be integrated into the error message, by utilizing %-formatting in Python. Although written in a different way compared to the developer patch, this patch achieves similar effect and passes the developer provided test-suite for this issue.
@@ 4. AI Program Improvement Framework
search_code (c) Search for code snippet c in the codebase. Region of code surrounding the searched snippet.
search_code_in_file (c, f) Search for code snippet c in file f. Region of code surrounding the searched snippet.
@@ 5. Experiment Setup
We selected two LLM-based agent systems Devin (Labs, 2024) and Swe-agent (Yang et al., 2024) as baselines and compare their performance against AutoCodeRover. Swe-agent is publicly available as a GitHub repository [^3] https://github.com/princeton-nlp/SWE-agent, so we replicated it as Swe-agent-rep in SWE-bench lite with the default setting based on the provided scripts. In contrast to Swe-agent, we do not have access to Devin, so we take the most relevant reported result from their technical report (Team, 2024). To avoid the natural randomness of LLM, we repeat our experiments with AutoCodeRover and Swe-agent-rep three times when possible, and report the average and total numbers across three repetitions. We use (1) the percentage of resolved instances, (2) average time cost, and (3) average token cost to evaluate the effectiveness of the tools. These evaluation metrics represent overall effectiveness, time efficiency, and economic efficacy in resolving real-world GitHub issues.
@@
We use the state-of-the-art OpenAI GPT-4 (gpt-4-0125-preview) as foundation inference model for both AutoCodeRover and Swe-agent-rep. In AutoCodeRover, the GPT-4 model is responsible for selecting search APIs to retrieve codebase context, refining the issue description, and writing a final patch. For parameters of GPT-4, we set a low temperature=0.2, max_tokens=1024 to produce relatively deterministic results and enable sufficient reasoning length for AutoCodeRover, and all other parameters remain as per default. For Swe-agent-rep, we increased the default cost limit of interacting with GPT-4 model (for resolving a task instace) to two USD and kept other settings as default. Note that, both two tools do not have a time limit, AutoCodeRover terminates either a patch is generated or the context retrieval stage repeats ten times. Swe-agent-rep terminates either a patch is generated or the cost of resolving an issue reaches two USD. Experiment results and artifacts for AutoCodeRover and Swe-agent-rep can be found at https://github.com/nus-apr/auto-code-rover.
@@
All experiments are conducted on an x86_64 Linux server with Ubuntu 20.04 installed.
@@ 6. Evaluation
6.1. RQ1: Overall Effectiveness on SWE-bench
We first measure the overall effectiveness of AutoCodeRover and baselines with the number of resolved task instances in SWE-bench. With the goal of understanding to what extent the current AI systems can automatically resolve real-life software issues, the only inputs we provided are the natural language issue description and a local code repository checked out at the erroneous version. We repeated the experiment of AutoCodeRover three times and presented the average and total number of resolved software issues across the three runs. The average and total results are denoted as ACR-avg and ACR-all respectively (for brevity, we use ACR to denote AutoCodeRover in this section). When reporting time and token/cost for ACR-all, we report the time and cost required for running each task three times. Since Devin was evaluated on a random 25% subset of SWE-bench (Labs, 2024), we also report results of AutoCodeRover on this subset (we refer to this as “SWE-bench Devin subset”). Table 2 shows the overall result in full SWE-bench, SWE-bench Devin subset, SWE-bench lite respectively. Figure 6 shows a visual summary of AutoCodeRover’s comparison with Swe-agent and Devin.
@@ 6.1. RQ1: Overall Effectiveness on SWE-bench
In the full SWE-bench, ACR-all (union of from the three ACR runs) resolved 15.95% task instances taking a time of 701 seconds per task (less than 12 minutes). In comparison, Swe-agent resolved 12.29% tasks in full SWE-bench, according to their report (Jimenez et al., 2024a) (we did not replicate Swe-agent-rep on the full SWE-bench due to the high cost).
@@
The state-of-the-art closed-source baseline tool Devin (Labs, 2024) is evaluated in a random 25% subset of SWE-bench. To compare AutoCodeRover with Devin, we report AutoCodeRover’s result on the 570 task instances Devin was evaluated on. The results of AutoCodeRover are taken from the ACR runs on full SWE-bench. In the SWE-bench Devin subset, the union of three runs of AutoCodeRover successfully resolved 15.79% of the task instances, which is higher than Devin. Figure 7(a) provides a more detailed exposition of the resolved tasks. Besides, the times taken by AutoCodeRover and Devin are comparable.
Results on SWE-bench lite.
We performed another round of experiments with AutoCodeRover on SWE-bench lite (300 instances). The results reported in Table 2 indicate that on average ACR-avg can resolve 16.11% task instances in SWE-bench lite, which is at par with the reported results from Swe-agent (Jimenez et al., 2024a). We also investigated the union of all resolved tasks in the three runs, in which the percentage of resolved task instances increased to 22.33%.
Table 2. Overall Result of AutoCodeRover (ACR) and baselines on full SWE-bench, SWE-bench Devin subset, and SWE-bench lite. ”-” indicates data not available.
@@
Swe-agent (Jimenez et al., 2024a) 12.29% (282) 93 -
ACR-avg 10.59% (243) 234 40177 (\$0.464)
ACR-all 15.95% (366) 701 120530 (\$1.392)
@@
Devin (Labs, 2024) 13.86% (79) ¿ 600 -
ACR-avg 10.47% (59.7) 231 39068 (\$0.452)
ACR-all 15.79% (90) 692 117204 (\$1.357)
Reported result on SWE-bench lite (size=300)
Swe-agent (Jimenez et al., 2024a) 17.00% (51) 93 69976 (\$0.739) [^4] This cell denotes the average cost per instance in our replication experiment (SWE-bench lite)
ACR-avg 16.11% (48.3) 173 37602 (\$0.435)
ACR-all 22.33% (67) 520 112806 (\$1.304)
Figure 6. Summary of Results
Detailed Comparison with Swe-agent
Since Swe-agent is publicly available, we also attempted to run Swe-agent on SWE-bench lite. We replicated Swe-agent with two USD as cost budget for conversation with LLM per task instance in our environment (denote as Swe-agent-rep, it terminates either a patch is generated or reaches the two USD budget). Table 3 shows that when considering the union of all resolved tasks across three repetitions, AutoCodeRover resolved 22.33% out of the 300 tasks, whereas Swe-agent-rep resolved 14.67%. We further analyzed the commonly and uniquely resolved instances between AutoCodeRover and Swe-agent-rep in Figure 7(b), and we found that AutoCodeRover and Swe-agent-rep complement each other in different scenarios. AutoCodeRover uniquely resolved 32 task instances, which benefited from the fine-grained code context search at the AST level to precisely locate the bug locations (e.g., django-13401 searches three necessary methods at one time). On the other hand, the main reason that AutoCodeRover failed on the 9 unique instances resolved by Swe-agent-rep, is unimplemented search APIs (e.g., search_file invoked in django-12286). In such cases, AutoCodeRover can generate invalid search results when unimplemented APIs are invoked. This implies more robust search APIs are desired for future improvement. Since the evaluation environment is important to the result, to make the experimental replication complete and thorough, we also evaluated the generated patches of 300 task instances of SWE-bench lite within the Swe-agent docker environment which has been publicly released. However, Table 3 shows that the number of resolved issues for both ACR and Swe-agent-rep decreased, indicating that the docker released by SWE-bench team may need more work.
@@
We are also interested in assessing the feasibility of deploying AutoCodeRover in the real world in terms of time and economic cost. On average, AutoCodeRover takes 234 seconds and 40177 tokens (equivalent to 0.464 USD) to resolve one task instance in SWE-bench. In comparison, our replication experiments with Swe-agent-rep cost 69976 tokens (equivalent to 0.739 USD) per task instance. When considering the combined three repetitions, AutoCodeRover takes 701 seconds (~11.7 minutes) per task, which is below the 30-60 minute time limit deemed acceptable by developers for automated repair tools (Noller et al., 2022). Looking into the 67 issues resolved by AutoCodeRover in SWE-bench lite, it costs on average ~2.77 days for developers to create pull requests for 57 issues, and the other 10 issues take even longer to be closed by developers (ranging from 34 - 4023 days). The short response time and low cost show the significant potential for AutoCodeRover to act as a first step in future autonomous bug fixing.
Table 3. The percentage of resolved task instances of AutoCodeRover and Swe-agent-rep on SWE-bench lite. The column “In Swe-agent Docker” represents the evaluation result under the docker environment released publicly by Swe-agent team on 6th April 2024.
Replication result on SWE-bench lite (size=300)
In our environment In Swe-agent Docker
ACR Agent-Rep ACR Agent-Rep
Run 1 16.00% (48) 9.33% (28) 10.00% (30) 6.67% (20)
Run 2 15.67% (47) 11.00% (33) 10.33% (31) 7.00% (21)
Run 3 16.67% (50) 9.33% (28) 10.67% (32) 6.00% (18)
All 22.33% (67) 14.67% (44) 14.00% (42) 9.00% (27)
(a) AutoCodeRover and Devin on SWE-bench Devin subset.
(b) AutoCodeRover and Swe-agent-rep on SWE-bench lite.
Figure 7. Venn diagrams of resolved tasks instances.
@@
Overfitting is a well-known challenge in the Automated Program Repair community (Gao et al., 2019). A program patch that passes the given test suite is said to be plausible. However, a plausible patch is deemed as overfitting if it fails to conform to the developer’s intent. Otherwise, it is deemed as correct. To further understand the patch quality of AutoCodeRover and baselines, we manually verify the correctness of task-resolving (i.e. plausible) patches in SWE-bench lite. Since three repetitions are performed, we consider a task to have a correct patch if any of the three repetitions produced a correct patch. A plausible patch is correct if it is semantically equivalent to the developer patch. In this verification process, at least two authors of the paper cross-validated each patch, and any disagreement was resolved with another author. Overall, on SWE-bench lite, ACR has a correctness rate of 65.7% (44 correct/67 plausible). Swe-agent-rep has a slightly higher correctness rate of 72.7% (32/44), but the absolute number of correctly resolved tasks is smaller than ACR. Finally, the correctness rate of Devin on SWE-bench Devin subset is 53.2% (42/79). We observed that vast majority of ACR’s overfitting patches (all but 2 of the overfitting patches) modify the same methods as the developer patches, but the code modifications are wrong. This means that even the overfitting patches from ACR are useful to the developer, since it helps in localization. The main causes of wrong modifications are the limits of the LLM’s capability or insufficient context. Apart from these, we noticed two other interesting causes of overfitting. One cause is that the issue creator gives a preliminary patch in the description. This patch can be different from the final developer patch, misleading the LLM. The other interesting cause is that the issue creator mentioned a case that needs to be handled. The LLM only fixes this mentioned case, but the developer fixed other similar cases as well. The two causes indicate that the issue description, just like test suites, can be an incomplete specification.

### v3

@@ Abstract.
Researchers have made significant progress in automating the software development process in the past decades. Automated techniques for issue summarization, bug reproduction, fault localization, and program repair have been built to ease the workload of developers. Recent progress in Large Language Models (LLMs) has significantly impacted the development process, where developers can use LLM-based programming assistants to achieve automated coding. Nevertheless, software engineering involves the process of program improvement apart from coding, specifically to enable software maintenance (e.g. program repair to fix bugs) and software evolution (e.g. feature additions). In this paper, we propose an automated approach for solving Github issues to autonomously achieve program improvement. In our approach called AutoCodeRover, LLMs are combined with sophisticated code search capabilities, ultimately leading to a program modification or patch. In contrast to recent LLM agent approaches from AI researchers and practitioners, our outlook is more software engineering oriented. We work on a program representation (abstract syntax tree) as opposed to viewing a software project as a mere collection of files. Our code search exploits the program structure in the form of classes/methods to enhance LLM’s understanding of the issue’s root cause, and effectively retrieve a context via iterative search. The use of spectrum-based fault localization using tests, further sharpens the context, as long as a test-suite is available. Experiments on the recently proposed SWE-bench-lite (300 real-life Github issues) show increased efficacy in solving Github issues (19% on SWE-bench-lite), which is higher than the efficacy of the recently reported Swe-agent. Interestingly, our approach resolved 57 GitHub issues in about 4 minutes each (pass@1), whereas developers spent more than 2.68 days on average. In addition, AutoCodeRover achieved this efficacy with significantly lower cost (on average, \$0.43 USD), compared to other baselines. We posit that our workflow enables autonomous software engineering, where, in future, auto-generated code from LLMs can be autonomously improved.
large language model, automatic program repair, autonomous software engineering, autonomous software improvement
copyright: rightsretaineddoi: 10.1145/3650212.3680384journalyear: 2024isbn: 979-8-4007-0612-7/24/09conference: Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis; September 16–20, 2024; Vienna, Austriabooktitle: Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA ’24), September 16–20, 2024, Vienna, Austriasubmissionid: issta24main-p1370-pccs: Software and its engineering Automatic programmingccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software testing and debuggingccs: Computing methodologies Natural language processing
@@ 1. Beyond Automatic Programming
Our contribution lies in the effective use of code search to make software engineering processes like program repair autonomous. Instead of viewing the codebase as a collection of files, we posit AutoCodeRover as an approach that gleans specification from software structure, which is used to guide the patching. The code search in AutoCodeRover is a vehicle for inferring specification from program structure. We demonstrate this capability by autonomously solving GitHub issues. We report favorable experimental results on the SWE-bench lite dataset (Jimenez et al., 2024b). We achieve 19% efficacy on the SWE-bench lite with 300 GitHub issues. Overall, we see the need to endow a software engineering oriented outlook to the recent flurry of activity on LLM agents for software engineering, which mostly has only an AI flavor. The angle of software engineering can be conveyed by the following five dimensions.
@@
  • We posit that higher efficacy of automated repair is more important than time efficiency, as long as the time is within a threshold. It is well-known in empirical software engineering research that time limits of 30-60 minutes for automated repair are tolerable, based on extensive developer surveys in the field (Noller et al., 2022). We thus report 19% efficacy on SWE-bench lite in solving GitHub issues, within 4 minutes. We can compare this time limit of 4 minutes to the average time of 2.68 days to fix the GitHub issues manually.
  • One should be able to exploit debugging techniques like test-based fault localization to guide the search for code in resolving GitHub issues. We show that the use of fault localization in setting the code context leads to an increase in the efficacy of AutoCodeRover in solving GitHub issues.
@@ 2.1. Program Repair
Recent work (Fan et al., 2023; Jiang et al., 2023; Xia et al., 2023) have shown the use of LLMs for automated program repair. This line of work often assumes the buggy program statements are given (i.e. perfect fault localization assumption) and focuses on constructing APR-specific prompts that guide LLM to generate a patch for the selected buggy program statements multiple times until a patch that passes all tests is found. The typical work targets repairing functional (Fan et al., 2023; Prenner et al., 2022; Wang et al., 2023), type (Chow et al., 2024) bugs, and even software vulnerabilities (Fu et al., 2022; Jin et al., 2023; Pearce et al., 2023a) in different languages with open-source/commercialised models (e.g., CodeT5, ChatGPT, GPT-4, etc). However, obtaining buggy locations for a large project is an essential and challenging task in resolving real-life bug reports.
@@
2.2. LLM Agents for SE and Dataset
SWE-bench lite (Jimenez et al., 2024b) is a benchmark that aims to evaluate the capabilities of large language models in resolving end-to-end real-life software engineering tasks. The benchmark consists of 300 real-life software engineering task instances collected from the repositories of 11 popular large Python projects  [^1] https://www.swebench.com/index.html (e.g., django, sympy). Each SWE-bench lite task instance contains a pair of Github issue and corresponding pull requests. The GitHub issue either reports a bug to be fixed or requests to implement new features. The pull request includes the code changes made by human developers to resolve the issue and test cases that prevent the issue. Unlike traditional code generation tasks in HumanEval (Chen et al., 2021) and MBPP (Li et al., 2022) benchmark, resolving a SWE-bench lite instance is particularly challenging because it requires automatically generating code changes that address the problem in a GitHub issue for a matured large code repository based only on the issue description. More specifically, the process may involve a series of complex tasks like reasoning the target bug location across files in the code repository, analyzing the root cause of the issue, proposing bug-fixing strategies, and eventually writing a patch that passes all the test cases added in the pull request. There are a few primary attempts at tackling tasks in SWE-bench lite. Devin (Labs, 2024) is named the first AI software engineer that can solve various software engineering tasks, including building a project from scratch, bug-fixing/feature-addition for existing projects. However, it is a close-sourced commercial tool, and its details are not available. Swe-agent (Yang et al., 2024) is a concurrent work against AutoCodeRover. It designed an agent-computer interface (ACI) that allows LLM agents to execute basic file operations via shell commands to achieve interaction between the LLM engine and a software repository. Compared to Swe-agent, we posit AutoCodeRover as an approach that gleans specification from software structure, which is used to guide the patching. Instead of viewing the codebase as a collection of files.
@@ 3. Motivating Example
""" AutoCodeRover patch. """
except (ValueError, TypeError, self.queryset.model.DoesNotExist):
- raise ValidationError(self.error_messages[’invalid_choice’],
- code=’invalid_choice’)
+ % Include the invalid value in the error message
+ raise ValidationError(
+ self.error_messages[’invalid_choice’] % {’value’: value},
+ code=’invalid_choice’)
""" Developer patch. """
except (ValueError, TypeError, self.queryset.model.DoesNotExist):
- raise ValidationError(self.error_messages[’invalid_choice’],
- code=’invalid_choice’)
+ raise ValidationError(
+ self.error_messages[’invalid_choice’],
+ code=’invalid_choice’,
+ params={’value’: value},)
@@
In Step 3, the agent receives the implementation of validate and to_python methods. Among the two methods, it selects to_python to be the more suitable place to make changes, since to_python raises the relevent exception and does not include the invalid value. At this point, the retrieval agent deems the collected code context as sufficient for understanding the issue and drafting the patch. The identified buggy location, together with the gathered context and analysis so far, is passed to another patch generation agent. This agent is instructed to write patches following the format described in Step 4 (see yellow box) of Figure 2. In Step 4, a patch is written to allow a value to be integrated into the error message, by utilizing %-formatting in Python. Figure 3 shows the patch generated by AutoCodeRover and the developer. Although written in a different way compared to the developer patch, this patch achieves similar effect and passes the developer provided test-suite for this issue.
@@ 4. AI Program Improvement Framework
search_code (c) Search for code snippet c in the codebase. +/- 3 lines of the searched snippet c.
search_code_in_file (c, f) Search for code snippet c in file f. +/- 3 lines of the searched snippet c.
@@ 5. Experiment Setup
We selected two LLM-based agent systems Devin (Labs, 2024), Swe-agent (Yang et al., 2024) as baselines and compare their performance against AutoCodeRover. We directly compare AutoCodeRover with Swe-agent’s original results as it is publicly available at GitHub [^3] https://github.com/princeton-nlp/SWE-agent. In contrast to Swe-agent, we do not have access to Devin, so we take the most relevant reported result from their technical report (Team, 2024). To avoid the natural randomness of LLM, we repeat our experiments three times. We report the result with the AutoCodeRover @1 and AutoCodeRover @3 annotations (i.e. pass@1, pass@3 metrics (Chen et al., 2021) respectively). We use (1) the percentage of resolved instances, (2) average time cost, and (3) average token cost to evaluate the effectiveness of the tools. These evaluation metrics represent overall effectiveness, time efficiency, and economic efficacy in resolving real-world GitHub issues.
@@
We use the state-of-the-art OpenAI GPT-4 (gpt-4-0125-preview) as foundation inference model for AutoCodeRover. In AutoCodeRover, the GPT-4 model is responsible for selecting search APIs to retrieve codebase context, refining the issue description, and writing a final patch. For parameters of GPT-4, we set a low temperature=0.2, max_tokens=1024 to produce relatively deterministic results and enable sufficient reasoning length for AutoCodeRover, and all other parameters remain as per default. AutoCodeRover terminates either a patch is generated or the context retrieval stage repeats ten times. Experiment results and artifacts for AutoCodeRover can be found at https://github.com/nus-apr/auto-code-rover.
@@
All experiments are conducted on an x86_64 Linux server with Ubuntu 20.04 installed. The correctness of generated patches by AutoCodeRover is evaluated on the official SWE-bench docker environment.
@@ 6. Evaluation
6.1. RQ1: Overall Effectiveness
We first measure the overall effectiveness of AutoCodeRover and baselines with the number of resolved task instances in SWE-bench. With the goal of understanding to what extent the current AI systems can automatically resolve real-life software issues, the only inputs we provided are the natural language issue description and a local code repository checked out at the erroneous version. We repeated the experiment of AutoCodeRover three times to avoid randomness and presented the results in pass@1 and pass@3 metrics (Chen et al., 2021), which are denoted as AutoCodeRover @1 and AutoCodeRover @3 respectively. When reporting time and token/cost for AutoCodeRover @3, we report the time and cost required for running each task three times. Since Devin was evaluated on a random 25% subset of SWE-bench (Labs, 2024), we also report results of AutoCodeRover on this subset (we refer to this as “SWE-bench Devin subset”). Table 2 shows the overall result of AutoCodeRover in SWE-bench lite, full SWE-bench and SWE-bench Devin subset and Figure 7 shows a visual efficacy summary of AutoCodeRover’s comparison with Swe-agent, Devin.
@@ 6.1. RQ1: Overall Effectiveness
The results reported in Table 2 indicate that in the full SWE-bench, AutoCodeRover @1 resolved 12.42% task instances with 248 seconds per task, and AutoCodeRover @3 resolved 17.96% task instances taking a time of 701 seconds per task, which is at par with 12.47% resolved tasks by the concurrent work Swe-agent, according to their technical report (Yang et al., 2024). We also performed another round of experiments with AutoCodeRover on SWE-bench lite (300 instances). AutoCodeRover @1 resolved 19% task instances in taking an average time of 195 seconds, whereas Swe-agent resolved 18% task instances. Moreover, we also investigated AutoCodeRover @3 in SWE-bench lite, in which the percentage of resolved task instances increased to 26%. The results of AutoCodeRover @3 in SWE-bench imply that AutoCodeRover may be complemented by running multiple times, highlighting the possibility of improving AutoCodeRover with an iterated generate-and-validate process in the future, if any program specification is available.
@@
To compare with Devin, we report AutoCodeRover’s result on the 570 task instances (random 25% subset of SWE-bench) Devin (Labs, 2024) was evaluated on. The results of Devin are taken from their technical report (Team, 2024). In this Devin particular subset, AutoCodeRover successfully resolved 12.63% task instances on average in pass@1, and 18.77% of the task instances in pass@3, which is higher than Devin. Besides, the time taken by AutoCodeRover is much smaller than Devin.
Detailed Comparison with Swe-agent
In the rest of RQ1, we compare AutoCodeRover with Swe-agent on SWE-bench lite using the highlighted AutoCodeRover @1 result. We further analyzed the commonly and uniquely resolved instances between AutoCodeRover and Swe-agent in Figure 6, and we found that AutoCodeRover and Swe-agent complement each other in different scenarios. AutoCodeRover uniquely resolved 26 task instances, which benefited from the fine-grained code context search at the AST level to precisely locate the bug locations (e.g., django-13401 searches three necessary methods at one time). On the other hand, the main reason that AutoCodeRover failed on the 23 unique instances resolved by Swe-agent is unimplemented search APIs (e.g., search_file invoked in django-12286). When such unimplemented APIs are invoked, AutoCodeRover returns an error message to the LLM, but the LLM may still fail to invoke valid search APIs in later attempts. This implies more robust search APIs are desired for future improvement of AutoCodeRover.
Table 2. Overall Result of AutoCodeRover and baselines on full SWE-bench, SWE-bench Devin subset, and SWE-bench lite. ”-” indicates data not available.
@@
Reported result on SWE-bench lite (size=300)
Swe-agent (Yang et al., 2024) 18.00% (54) - 245k (\$2.51)
AutoCodeRover @1 19.00% (57) 195 37k (\$0.43)
AutoCodeRover @3 26.00% (78) 520 112k (\$1.30)
ACR-sbfl 22.00% (66) 250 40k (\$0.47)
@@
Swe-agent (Yang et al., 2024) 12.47% (286) - 240k (\$2.46)
AutoCodeRover @1 12.42% (285) 248 39k (\$0.45)
AutoCodeRover @3 17.96% (422) 701 120k (\$1.39)
@@
Swe-agent (Yang et al., 2024) 13.51% (77) - 234k (\$2.40)
Devin (Labs, 2024; Team, 2024) [^4] The reported result of Devin is evaluated on a random 25% subset of full SWE-bench 13.86% (79) ¿ 600 -
AutoCodeRover @1 12.63% (72) 238 37k (\$0.42)
AutoCodeRover @3 18.77% (107) 692 117k (\$1.36)
Figure 6. Venn diagrams of resolved tasks instances by AutoCodeRover and Swe-agent, on SWE-bench lite.
Figure 7. Task-resolving success rates (%) and average costs per task in USD of AutoCodeRover and baselines.
Figure 8. The number of resolved tasks and token cost distributions of AutoCodeRover and Swe-agent.
@@
Today, the price of invoking state-of-the-art LLMs such as GPT-4 and Claude-3-Opus are still very expensive. Hence, we are also interested in assessing the feasibility of deploying AutoCodeRover in the real world in terms of time and economic cost. Figure 7 shows the comparison between task-resolving success rate and average costs per task for all tools. On average, AutoCodeRover takes 195 seconds and  37k tokens (equivalent to 0.43 USD) to resolve one task instance in SWE-bench lite. In comparison, Swe-agent costs 245k tokens (equivalent to 2.51 USD) per task instance. The cost of Devin is empty because it is not publicly available. When considering the combined three repetitions, AutoCodeRover takes 520 seconds (~8.67 minutes) per task, which is below the 30-60 minute time limit deemed acceptable by developers for automated repair tools (Noller et al., 2022). Looking into the 78 issues resolved by AutoCodeRover @3 in SWE-bench lite, it costs on average ~2.68 days for developers to create pull requests for 66 issues, and the other 12 issues take even longer to be closed by developers (ranging from 34 - 4023 days). The short response time and significantly low success-rate/cost ratio show the potential for AutoCodeRover to act as a first affordable step in future autonomous bug fixing.
Token Distribution of Resolved Tasks
We dived into the details of token cost distribution among resolved task instances by AutoCodeRover and Swe-agent in Figure 8. The x-axis represents token cost ranges and the y-axis represents the number of resolved tasks. Figure 8 shows that regarding the resolved tasks, AutoCodeRover is even more token-efficient. AutoCodeRover resolved 19 tasks with less than 10k tokens (~0.1 USD) and 31 tasks between 10k-30k tokens, whereas 66.7% of resolved tasks by Swe-agent at least requires more than 100k tokens. This implies the strong capability of stratified context retrieval in AutoCodeRover to fast pinpoint error locations and thus produce final patch.
@@
Overfitting is a well-known challenge in the Automated Program Repair community (Gao et al., 2019). A program patch that passes the given test suite is said to be plausible. However, a plausible patch is deemed as overfitting if it fails to conform to the developer’s intent. Otherwise, it is deemed as correct. To further understand the patch quality of AutoCodeRover and baselines, we manually verify the correctness of task-resolving (i.e. plausible) patches in SWE-bench lite. Since three repetitions are performed, we consider a task to have a correct patch if any of the three repetitions produced a correct patch. A plausible patch is correct if it is semantically equivalent to the developer patch. In this verification process, at least two authors of the paper cross-validated each patch, and any disagreement was resolved with another author. Overall, on SWE-bench lite, AutoCodeRover has a correctness rate of 65.4% (51 correct/78 plausible). We observed that the vast majority of AutoCodeRover’s overfitting patches (all but 2 of the overfitting patches) modify the same methods as the developer patches, but the code modifications are wrong. This means that even the overfitting patches from AutoCodeRover are useful to the developer, since it helps in localization. The main causes of wrong modifications are the limits of the LLM’s capability or insufficient context. Apart from these, we noticed two other interesting causes of overfitting. One cause is that the issue creator gives a preliminary patch in the description. This patch can be different from the final developer patch, misleading the LLM. The other interesting cause is that the issue creator mentioned a case that needs to be handled. The LLM only fixes this mentioned case, but the developer fixed other similar cases as well. The two causes indicate that the issue description, just like test suites, can be an incomplete specification.